A corporate jet costs $1,350,000 and will incur $200,000 per year in fixed cost (maintenance, …) and $277 per hour variable cost (fuel, …). The jet will be operated 1200 hours per year for 5 years and then sold for $650,000. The jet revenues $1,000 per hour. The MARR is 15% per year.

Determine the Equivalent Uniform Annual Cost (EUAC) of the jet.
Answer #1

;Answer:

Jet cost J=1350000

Annual Fixed cost F=200000

Variable cost per Hour V=277

Revenue per hour R=1000

Number of Hours in a Year N=1200

So Total net revenue per year A=(R-V)*N=(1000-277)*1200=867600

Salavage value S= 650000

MARR r=15%

Number of Years n=5

EUAC= Equivalent annual cost of Jet for 5 years + Fixed cost per year -Total Net Reveneu per year - Equivalent annual of salvage value for 5 years

EUAC=J*r/(1-(1+r)^-n) + F-A-S*r/((1+r)^-n-1)

EUAC=1350000*15%/(1-(1+15%)^-5) +200000-867600-650000*15%/((1+15%)^-5-1)

EAUC =-$361279.11

Negative sign implies that we are making profit each year.